OpenStudy (anonymous):

A plane is flying due east with the velocity of 90m/s. The wind is blowing out of the north at 4m/s. What is the magnitude of the plane's resultant velocity? (Round answer to the nearest tenth.) Whoever answers, could you draw it out in your explanation please?

OpenStudy (zehanz):

See image. Using the Pythagorean Theorem: \[v_r=\sqrt{90^2+4^2}=\sqrt{106} \approx 10.3 m/s\]

OpenStudy (anonymous):



OpenStudy (anonymous):

Well . . . we're using The Triangle Method or the Tail-to-Tail Method.

OpenStudy (zehanz):

Both methods have been used by MeiTU and myself...

OpenStudy (anonymous):

The way that your diagrams show and the words that you used are nothing like my notes.

OpenStudy (zehanz):

Could you show your notes, then?

OpenStudy (anonymous):

I can try... An airplane's velocity through the air is a combo of its speed & its direction of flight. Velocity can be represented by a directed line segment called a vector. The length of the vector represents that magnitude, in this case, speed, while the arrowhead indicates the direction in which this magnitude is acting.  Then my notes say there are 2 ways vectors A&W can be added geometrically, the tail-to-tail and the triangle/tip-to-tail methods. The Tail-to-Tail Method notes say to redraw the 2 vector arrows so that the tails of each other are touching, while mkaing sure that the length and direction of each remain the same. Next, draw parallel vectors from the heads of each original vector. (This creates a parallelogram.) Now draw a 3rd vector from the tails of the original vectors to their heads. The new vector, r, is the sum of vectors A&W. Otherwise known as the resultant.